# Network Intrusion Datasets: A Survey, Limitations, and Recommendations

### v1

@@ 5.2.1 Intrusion-Specific
DoS/DDoS: The most common intrusion-specific datasets focus on Denial of service (DoS) and Distributed DoS (DDoS). These attacks have been around since the beginning of the Internet and still remain a major threat nowadays [170]. The first DoS dataset, LLDOS [90] (2000), was published by DARPA. Rare and valuable real-world samples were later published in CAIDA DDoS Attack 2007 [85] and Booters [53] datasets.
@@
Seeking to address challenges in obtaining real-world traffic, DDoS 2016 [100] was generated synthetically with selected per-packet features. Similar per-packet features are also present in Boğaziçi University DDoS Dataset [129], mixing emulated DDoS attacks with real campus traffic. CIC-DDoS2019 [119] emulates diverse flooding attacks in both raw packets and CICFlowMeter features. In contrast to flooding attacks, the 2017-SUEE-data-set [84] and USB-IDS-1 [145] focused on layer 7 slow DoS attacks, with the latter evaluating their effectivity against in-built webserver defensive modules.
@@
Botnet: Botnet detection is another crucial branch of NID research. The first dataset, ISOT Botnet [94] (2004-2010), mixes Storm and Walowdac botnet captures with non-malicious traffic from the Ericsson Research Lab in Hungary. A comprehensive set of 13 real-world botnet scenarios is provided in CTU-13 [55] by running unrestricted malware in a sandbox environment within an academic network. Selected CTU-13 traces and legitimate ISCX-IDS-2012 traffic were combined into the HogZilla [95] dataset with custom flow-based features. ISCX-IDS-2012 and CTU-13 data were also used with the ISOT Botnet and merged into the Botnet [96] dataset. In contrast to detection from entire traffic, ISOT HTTP Botnet [110] focuses on botnet detection only via DNS data. Botnet detection is also deeply intertwined with IoT, as IoT devices were historically poorly secured, resulting in their ease of compromise and consequent participation in a botnet. We cover them in the environment-specific datasets subsection.
@@
Anomaly detection: Instead of identifying concrete intrusions, anomaly detectors learn patterns of regular communication and then detect deviations from them. While any dataset can be used to benchmark anomaly-based systems, some were designed especially for this purpose. Comprehensive, Multi-Source Cyber-Security Events [101], referred to as Kent 2016 and Unified Host and Network [105] datasets, collected from the Los Alamos National Laboratory enterprise network, provide flow and end-host log data with a minimalist feature set. In addition, they are heavily anonymized and provide no direct labeling, significantly limiting their usability.
@@
MaWILab [46], the longest-running intrusion detection capture (ongoing since 2001), utilizes the MAWI Working Group Traffic Archive [171] for daily 15-minute backbone traces. These traces are automatically labeled via combined anomaly detectors. However, the project’s inactivity since 2015 likely renders detectors outdated, hurting the ground truth accuracy. Furthermore, despite the MAWI Archive continuing to provide daily traces, MAWILab includes limited recent data.
@@
Automatic labeling, this time based on external third-party Cyber Threat Intelligence, was also utilized for LuFlow [147]. It provides flow-based traffic from the Lancaster University network, labeling its data as benign, malicious, and outlier. The TrabID dataset [31] aims to realistically evaluate NIDSs’ adaptability by dividing the data into three types likely encountered in operational networks: known, similar, and unknown. Unfortunately, the environment details were never released.
@@
APT: Regular IDS datasets usually focus on a certain set of attacks performed in a non-systematic way. In contrast, Advanced Persistent Threat (APT) datasets simulate real-world cyber incidents, which typically follow a well-defined sequence of steps, as outlined in the MITRE ATT&CK Matrix [167]. APT datasets aim to mimic such a realistic APT behavior, typically starting with reconnaissance and ending with attacker’s objectives like data exfiltration, while often trying to keep such activities stealthy.
@@
DAPT 2020 [130] captures all APT stages performed by a red team (experienced cyber-attackers) in a virtualized environment. Realism was further increased by heavily unbalancing the traffic and including stealthy attacks, causing baseline semi-supervised detectors to perform poorly. The Unraveled [144] dataset expands on this work with a more complex network, longer duration, defender’s responses, attacks’ tracks covering, and different attacker types. Professional red teamers were also used in the PWNJUTSU [143] dataset, which is, unfortunately, unlabeled. Another APT activity, spanning 2.5 hours, is also simulated by the CyberFORCE Scenario [139]. Unfortunately, its data are labeled indirectly, and the corresponding documentation lacks many details.
@@
Others: In addition to the mentioned datasets, several others focused on specific intrusions are also available. Homoliak et al. [92] published ANSM-CDX, ANSM-TUN, and ANSM-NPBO for detecting buffer overflow attacks using custom flow-based features. SSHCure [97] dataset captures real-world SSH brute-force attempts via a university honeypot. SR-BH 2020 [134] offers flow-based data for evaluating HTTP(S) multi-label intrusion detection models. In contrast to standalone datasets, NDSec-1 [102] is designed for salting by injecting its traffic into other datasets. Finally, HIKARI-2021 [141] promotes the usage of encrypted traffic for NID purposes by employing application-layer attacks via HTTPS, extracting its features using CICFlowMeter and Zeek.
@@ 5.2.2 Environment-Specific
Unlike datasets captured in traditional computer networks, those described in the following paragraphs originate from more specialized environments. These environments include specialized network infrastructures like Software Defined Networking (SDNs) or end-host devices like the Internet of Things (IoT). As a result, they produce unique benign traffic patterns and are susceptible to environment-specific attacks, requiring tailored intrusion detection systems.
IoT (botnet): The Internet of Things (IoT) is the most prevalent specialized environment among the public NID datasets. These datasets incorporate IoT devices such as sensors, cameras, and smart home appliances, often connected to traditional networks via Wi-Fi (IEEE 802.11) [172] or specific technologies like ZigBee (IEEE 802.15.4) [173]. Historically, IoT devices were manufactured without cybersecurity considerations, causing IoT malware infections and consequent participation in botnets. Reflecting this trend, datasets IoT host-based ID dataset [111], N-BaIoT [117], IoT-23 [118], and MedBIoT [125] focus on botnet activity. Captured in controlled environments, these datasets simulate a botnet by infecting testbed hosts with popular malware families like Mirai and monitoring their behavior. In addition to botnet activity, the IoT network intrusion dataset [121] includes other attacks like scanning and ARP spoofing. Since this dataset provides only raw traffic, Ullah et al. applied the CICFlowMeter tool to it and produced the IoTID [122] dataset.
@@
CCD-INID-V1 [138] captures DoS, man-the-middle, and brute-force attacks within a network of four Raspberry Pi devices, using NFStream for flow-based feature extraction. Traffic from 60 physical IoT devices was assembled within the CIC IoT dataset 2022 [146]. Although primarily for monitoring, flood and brute-force attacks were added for intrusion detection, providing the data in raw format and via custom per-packet and statistical features. CICIoT2023 extends its scope to 105 devices and 33 attacks across seven categories, including the Mirai botnet. Along with the raw data, pre-extracted fixed-size packet window features are also available.
@@
IIoT/ICS: Industrial IoT (IIoT) creates an ecosystem of devices, sensors, and applications that collect, monitor, and analyze data from industrial operations, oftentimes integrating with Industrial Control Systems (ICS) technologies, e.g., Supervisory Control and Data Acquisition (SCADA). IIoT datasets aim to capture the behavior of industrial networks but also contain regular, non-industrial-specific attacks. WUSTL-IIoT-2018 [174] and WUSTL-IIOT-2021 [126] present reconnaissance and exploit attacks (the latter also DoS and command injection) within a small SCADA environment with a water storage tank. Their features were extracted by Argus in a bi-flow format. X-IIoTID [128] designed connectivity- and device-agnostic emulation in a physical testbed for various attacks categorized by the ATT&CK framework. Edge-IIoTset [140] proposed a small but diverse multi-layer testbed with numerous device types. It provides a rich set of attacks, including man-in-the-middle, malware, injection, information gathering, and DoS, offering both raw and pre-extracted data.
@@
ICS: While previous datasets mixed both IIoT and ICS traffic, ICS-Flow [150] focuses solely on the ICS environment, simulating a bottle-filling factory within a virtual testbed. Four attack types: reconnaissance, replay, DDoS, and man-in-the-middle were performed, and their activity was captured via ICSFlowGenerator as custom flow-based features. Using the same setup as ICS-Flow, the Westermo [158] dataset added anomaly detection scenarios like IP address misconfiguration. Since three different capture points were used, the data is also suitable for federated learning. ROSIDS23 [157] provides attacks against the Robot Operating System (ROS), utilizing CICFlowMeter for feature extraction. Finally, the OPC UA [132] dataset captures data from an OPC UA cyber-physical system with DoS, man-in-the-middle, and spoofing attacks.
@@
Other: Various other datasets cover additional environments. InSDN [131], SDN Dataset [133], and UNR-IDD [76] present regular, as well as environment-specific attacks for Software Defined Networking (SDN). Wireless-specific attacks like disassociation and deauthentification are presented by AWID2 [64] and AWID3 [137], emulating small-office-home-office (SOHO) networks. ISOT-CID [113] offers cloud-specific traces with network and end-host logs collected from three hypervisors in a production environment. Finally, the VOIP Enterprise – Attack Dataset [135] focuses on Voice over IP (VoIP)-specific attacks like VoIP-DoS.
@@ 5.3 Datasets Beyond This Survey
For the purpose of this survey, we define explicit, non-ambiguous criteria to filter relevant NID datasets. This caused the exclusion of several datasets commonly cited in other works (e.g., LBNL [175] or ADFA [176]). Additionally, we excluded dozens of relevant datasets due to their public unavailability. As these datasets might suit specific use cases, we list all exclusions along with their reasons in our GitHub repository (see Section 1).
Although network intrusion data backed with scientific publications (the focus of this survey) are a preferred way for benchmarking NIDSs, other sources might also provide relevant datasets. This subsection outlines these alternatives.
@@ 5.3.1 Data Repositories
Data repositories are centralized locations to store, share, and categorize the data. For cybersecurity and intrusion detection purposes, the most relevant include IEEE DataPort [^5] IEEE DataPort: https://ieee-dataport.org/, Kaggle [^6] Kaggle: https://www.kaggle.com/, and Zenodo [^7] Zenodo: https://zenodo.org/, hosting many datasets from scientific publications. Kaggle and Zenodo offer open access, while some data in DataPort might require a paid subscription. Additional general-purpose repositories are listed on Wikipedia [^8] Wikipedia datasets list: https://en.wikipedia.org/wiki/List_of_datasets_for_machine-learning_research.
In addition to general-purpose repositories with unrestricted upload, several repositories focused on network and malware traces also exist. These are typically maintained by an individual or a group and thus have their upload restricted. We list some of the most relevant below:
  • ANT Datasets [^9] ANT datasets: https://ant.isi.edu/datasets/index.html – maintained by the ANT lab at the University of Southern California. Contains various benign and malicious captures since 2006.
  • AZSecure-data [^10] AZSecure-data: https://www.azsecure-data.org/home.html – a collection of various cybersecurity datasets including several NID datasets as well.
@@
  • Contagio [^12] Contagio: https://contagiodump.blogspot.com/ – a blog with historic and contemporary malware samples, threats, observations, and traffic captures from emulated environments.
@@
  • Security Datasets [^15] Security Datasets: https://securitydatasets.com/i – an open-source project with dozens of malicious and benign datasets from various platforms.
@@ 5.3.2 Data Lists
One such list of publicly available PCAP files, primarily from malware activity, is maintained by NETRESEC [^18] NETRESEC list: https://www.netresec.com/?page=PcapFiles. On the other hand, researcher Laurens D’Hooge provides pre-processed NID datasets in a textual format [^19] L. D’Hooge data: https://www.kaggle.com/dhoogla/datasets. Several network and malware captures are also aggregated by the SecRepo project [^20] SecRepo: https://www.secrepo.com/. The Outlier Detection Datasets (ODDS) project [^21] ODDS: https://odds.cs.stonybrook.edu/ focuses on datasets for anomaly detection, listing a few cybersecurity ones as well. Finally, GitHub pages Awesome Cybersecurity Datasets [^22] Awesome Cybersecurity Datasts: https://github.com/shramos/Awesome-Cybersecurity-Datasets and Real-CyberSecurity-Datasets [^23] Real-CyberSecurity-Datasets: https://github.com/gfek/Real-CyberSecurity-Datasets also compile numerous cybersecurity datasets, including NID.

### v2

@@ 5.2.1 Intrusion-Specific
DoS/DDoS: Denial of Service (DoS) and Distributed DoS (DDoS) are the among the most prevalent intrusion-specific datasets. These attacks have been around since the beginning of the Internet and remain a major threat to this day [171]. The first DoS dataset, LLDOS [91] (2000), was published by DARPA. Rare and valuable real-world samples were published in CAIDA DDoS Attack 2007 [86] and Booters [57] datasets.
@@
Seeking to address challenges in obtaining real-world traffic, DDoS 2016 [101] was emulated within a virtualized environment, providing selected per-packet features. Similar per-packet features are also present in Boğaziçi University DDoS Dataset [131], mixing emulated DDoS attacks with real campus traffic. CIC-DDoS2019 [120] emulates diverse flooding attacks in raw packets and CICFlowMeter features. In contrast to flooding attacks, the 2017-SUEE-data-set [85] and USB-IDS-1 [145] focus on layer 7 slow DoS, with the latter evaluating its effectiveness against in-built webserver defensive modules.
@@
Botnet: Botnet detection is another crucial branch of NID research. The first dataset, ISOT Botnet [95] (2004-2010), mixes Storm and Walowdac botnet captures with non-malicious traffic from the Ericsson Research Lab in Hungary. A comprehensive set of 13 real-world botnet scenarios is provided in CTU-13 [59] by running unrestricted malware in a sandbox environment within an academic network. Selected CTU-13 traces and legitimate ISCX-IDS-2012 traffic were combined into the HogZilla [96] dataset with custom flow-based features. ISCX-IDS-2012 and CTU-13 data were also used with the ISOT Botnet and merged into the Botnet [97] dataset. In contrast to analyzing entire network traffic, ISOT HTTP Botnet [111] focuses on detecting botnets only via DNS data. Botnet detection is also closely linked to IoT, as IoT devices were historically poorly secured, making them easy targets for compromise and subsequent participation in botnets. We cover them in the environment-specific datasets subsection.
@@
Anomaly detection: Instead of identifying specific intrusions, anomaly detectors learn patterns of regular communication and then detect deviations from them. While any dataset can be used to benchmark anomaly-based systems, some were specifically designed for this purpose. Comprehensive, Multi-Source Cyber-Security Events (known as Kent 2016[102] and Unified Host and Network [106] datasets, collected from the Los Alamos National Laboratory enterprise network, provide flow and end-host log data with a base minimalistic feature set. In addition, heavy anonymization and the absence of direct labeling significantly limit their usability.
@@
MaWILab [50], the longest-running intrusion detection capture (since 2001), utilizes the MAWI Working Group Traffic Archive [172] for daily 15-minute backbone traces. These traces are automatically labeled via combined anomaly detectors. However, the project’s inactivity since 2015 has likely rendered these detectors outdated, hurting the ground truth accuracy. Furthermore, although the MAWI Archive continues to provide daily traces, MAWILab includes limited recent data.
@@
Automatic labeling, this time based on external third-party Cyber Threat Intelligence, was also utilized by LuFlow [148]. It provides flow-based traffic from the Lancaster University network, labeling its data as benign, malicious, or outlier. The TrabID dataset [32] aims to realistically evaluate NIDSs’ adaptability by dividing the data into three types likely encountered in operational networks: known, similar, and unknown. Unfortunately, its environment details were never released.
@@
APT: Conventional IDS datasets typically focus on a certain set of attacks executed in a non-systematic way. In contrast, Advanced Persistent Threat (APT) datasets simulate real-world cyber incidents, which typically follow a well-defined sequence of steps, as outlined in the MITRE ATT&CK Matrix [168]. APT datasets aim to replicate realistic APT behavior, typically starting with reconnaissance and progressing toward attacker’s objectives like data exfiltration, while often trying to keep such activities stealthy.
@@
DAPT 2020 [133] captures all APT stages performed by a red team (experienced cyber-attackers) in a virtualized environment. Realism was further increased by heavily unbalancing the traffic and including stealthy attacks, causing baseline semi-supervised detectors to perform poorly. The Unraveled [144] dataset expands on this work with a more complex network, longer duration, defender’s responses, attack tracks covering, and different attacker types. Professional red teamers were also used in the PWNJUTSU [143] dataset, which is, unfortunately, unlabeled. Another APT activity, spanning 2.5 hours, is also simulated by the CyberFORCE Scenario [140]. Unfortunately, its data are labeled indirectly, and the corresponding documentation lacks many details.
@@
Others: In addition to the mentioned datasets, several others focused on specific intrusions are also available. Homoliak et al. [93] published ANSM-CDX, ANSM-TUN, and ANSM-NPBO for detecting buffer overflow attacks using custom flow-based features. The SSHCure [98] dataset captures real-world SSH brute-force attempts via a university honeypot. SR-BH 2020 [136] offers flow-based data for evaluating HTTP(S) multi-label intrusion detection models. In contrast to standalone datasets, NDSec-1 [103] is designed for salting by injecting its traffic into other datasets. Finally, HIKARI-2021 [147] promotes the usage of encrypted traffic for NID purposes by employing application-layer attacks via HTTPS, extracting its features using CICFlowMeter and Zeek.
@@ 5.2.2 Environment-Specific
Unlike datasets collected from traditional computer networks, the following paragraphs focus on those originating from specialized environments. These include network infrastructures such as Software Defined Networking (SDNs) or the Internet of Things (IoT). As a result, they produce unique benign traffic patterns and are susceptible to environment-specific attacks, requiring tailored intrusion detection systems.
IoT (botnet): The Internet of Things (IoT) is the most prevalent specialized environment among the public NID datasets, featuring IoT devices such as sensors, cameras, or smart home appliances. Often connected to traditional networks via Wi-Fi (IEEE 802.11) [173] or IoT technologies like ZigBee (IEEE 802.15.4) [174], these devices were manufactured without cybersecurity considerations. Some are thus vulnerable to IoT malware infections and consequent participation in botnets.
Reflecting this trend, IoT host-based ID dataset [112], N-BaIoT [118], IoT-23 [119], and MedBIoT [126] focus on botnet activity. Captured in controlled environments, these datasets simulate a botnet by infecting testbed hosts with popular malware families like Mirai and monitoring their behavior. In addition to botnet activity, the IoT network intrusion dataset [122] includes other attacks like scanning and ARP spoofing. Since this dataset provides only raw traffic, Ullah et al. processed its data using CICFlowMeter, producing the IoTID [123] dataset.
@@
CCD-INID-V1 [132] captures DoS, man-the-middle, and brute-force attacks within a network of four Raspberry Pi devices, using NFStream for flow-based feature extraction. Traffic from 60 physical IoT devices is gathered within the CIC IoT dataset 2022 [146]. Although primarily designed for monitoring purposes, it also includes flood and brute-force intrusion scenarios. Its data are provided as raw PCAP and custom per-packet statistical features. CICIoT2023 [154] extends its scope to 105 devices and 33 attacks across seven categories, including the Mirai botnet. Along with the raw data, pre-extracted fixed-size packet window features are also available.
IIoT/ICS: Industrial IoT (IIoT) is an ecosystem of devices, sensors, and applications that collect, monitor, and analyze data from industrial operations [175], oftentimes integrating Industrial Control Systems (ICS) technologies, e.g., Supervisory Control and Data Acquisition (SCADA). IIoT datasets aim to capture the behavior of industrial networks but some also contain regular, non-industrial-specific attacks.
@@
WUSTL-IIoT-2018 [176] and WUSTL-IIOT-2021 [127] present reconnaissance and exploit attacks (the latter also DoS and command injection) within a small SCADA environment with a water storage tank. Their features were extracted by Argus in a bi-flow format. X-IIoTID [130] designed connectivity- and device-agnostic emulation in a physical testbed for various attacks categorized by the ATT&CK framework. Edge-IIoTset [141] proposed a small but diverse multi-layer testbed with numerous device types. It provides a rich set of attacks, including man-in-the-middle, malware, injection, information gathering, and DoS, offering both raw and pre-extracted data.
@@
ICS: While previous datasets mixed both IIoT and ICS traffic, ICS-Flow [151] focuses solely on the ICS environment, simulating a bottle-filling factory within a virtual testbed. Four attack types: reconnaissance, replay, DDoS, and man-in-the-middle were executed, and their activity was captured via ICSFlowGenerator as custom flow-based features. Using the same setup as ICS-Flow, the Westermo [159] dataset added anomaly detection scenarios like IP address misconfiguration. Since three different capture points were used, the data is also suitable for federated learning. ROSIDS23 [158] provides attacks against the Robot Operating System (ROS), utilizing CICFlowMeter for feature extraction. Finally, the OPC UA [134] dataset captures data from an OPC UA cyber-physical system with DoS, man-in-the-middle, and spoofing attacks.
@@
Other: Various other datasets cover additional environments. InSDN [129], SDN Dataset [135], and UNR-IDD [77] present regular, as well as environment-specific attacks for Software Defined Networking (SDN). Wireless-specific attacks like disassociation and deauthentification are captured by AWID2 [65] and AWID3 [139], emulating small-office-home-office (SOHO) networks. ISOT-CID [114] offers cloud-specific traces with network and end-host logs collected from three hypervisors in a production environment. Finally, the VOIP Enterprise – Attack Dataset [138] focuses on Voice over IP (VoIP)-specific attacks like VoIP-DoS.
@@ 5.3 Datasets Beyond This Survey
For this survey, we established explicit, non-ambiguous criteria to filter relevant works, leading to the exclusion of several well-known datasets (e.g., LBNL [177] or ADFA [178]). Additionally, we excluded dozens of relevant datasets due to being publicly unavailable. As these datasets might be suitable for specific use cases, we provide a complete list of exclusions and their reasons on our GitHub repository (see Section 1).
Although network intrusion datasets backed by scientific publications (the focus of this survey) are a preferred way for benchmarking NIDSs, other sources might also provide relevant data. This subsection outlines these alternatives.
@@ 5.3.1 Data Repositories
Data repositories are centralized locations to store, share, and categorize the data. For cybersecurity and intrusion detection purposes, the most relevant include IEEE DataPort [^5] IEEE DataPort: https://ieee-dataport.org/, Kaggle [^6] Kaggle: https://www.kaggle.com/, and Zenodo [^7] Zenodo: https://zenodo.org/, hosting many datasets from scientific publications. Kaggle and Zenodo offer open access, while data on DataPort might require a paid subscription. Additional general-purpose repositories are listed on Wikipedia [^8] Wikipedia datasets list: https://en.wikipedia.org/wiki/List_of_datasets_for_machine-learning_research.
In addition to general-purpose repositories with unrestricted upload, several repositories specifically focused on network and malware traces also exist. These are typically maintained by individuals or groups, and therefore, their uploads are often restricted. We list some of the most relevant below:
  • ANT Datasets [^9] ANT datasets: https://ant.isi.edu/datasets/index.html – maintained by the ANT lab at the University of Southern California. It contains various benign and malicious captures since 2006.
  • AZSecure-data [^10] AZSecure-data: https://www.azsecure-data.org/home.html – a collection of various cybersecurity datasets, including several NID datasets as well.
@@
  • Contagio [^12] Contagio: https://contagiodump.blogspot.com/ – a blog with historical and contemporary malware samples, threats, observations, and traffic captures from emulated environments.
@@
  • Security Datasets [^15] Security Datasets: https://securitydatasets.com/ – an open-source project with dozens of malicious and benign datasets from various platforms.
@@ 5.3.2 Data Lists
A list of publicly available PCAP files, primarily sourced from malware activity, is maintained by NETRESEC [^18] NETRESEC list: https://www.netresec.com/?page=PcapFiles. In contrast, researcher Laurens D’Hooge focuses on pre-processed NID datasets in a textual format [^19] L. D’Hooge data: https://www.kaggle.com/dhoogla/datasets. Several network and malware captures are also aggregated by the SecRepo project [^20] SecRepo: https://www.secrepo.com/. The Outlier Detection Datasets (ODDS) project [^21] ODDS: https://odds.cs.stonybrook.edu/ collects datasets for anomaly detection, including some for cybersecurity. Finally, the GitHub repositories Awesome Cybersecurity Datasets [^22] Awesome Cybersecurity Datasts: https://github.com/shramos/Awesome-Cybersecurity-Datasets and Real-CyberSecurity-Datasets [^23] Real-CyberSecurity-Datasets: https://github.com/gfek/Real-CyberSecurity-Datasets compile numerous cybersecurity datasets, including those for NID.